# Computational Modelling of Thixotropic Multiphase Fluids

### v1

@@ 4.3 Droplet dynamics in a periodically constricted channel
To further investigate the behavior of thixotropic fluids, we analyze the transition of a droplet through a periodically constricted channel. The experiment focuses on observing the changes in viscosity during the contraction and subsequent relaxation of the droplet as it moves through the chamber, driven by the abrupt variations in shear rate between different regions of the domain. To correctly analyze the thixotropic phenomenon it is necessary to define the time scales. As indicated by Vázquez-Quesada et al [60], the sonic time $\tau_{s}=R/c$ which is referred to the time in which a sound wave propagates over a distance $R$ at a sound velocity $c$, must be less than the viscous time ($\tau_{visc}$) to satisfy incompressibility. To determine the viscous relaxation time of the system, we analyze the time decay of the characteristic velocity of the fluid. Considering that in processes dominated by viscous dissipation the velocity follows a decreasing exponential behavior of the form $v(t)\propto\exp(-t/\tau_{visc})$, we calculate $\ln(v)$ as a function of time, thus obtaining a linear relationship. From the linear fit of these data, we relate the slope $m$ of the line to the relaxation time as $\tau_{visc}=-1/m$. These two time scales must meet the condition:
[…]
where $T=L/U$ is the fluid characteristic time. The characteristic thixotropic time scale $\lambda_{0}$ is chosen here on the same order of the characteristic flow time $T$, i.e., $\tau_{s}\ <<\ \tau_{visc}\ <<\ \lambda_{0},\ T$.
We employ for the simulations a channel with periodical boundary conditions of total length 32. Each of the two cross sections has length $L=16$ and a height of the constrained section $H_{c}=2.2$ and for the relaxed section $H_{r}=10$. We consider the transporting fluid to be Newtonian with viscosity $\eta=1$. We treat the droplet as thixotropic with properties $\eta_{\infty}=10$ and $\alpha=4$. We vary the thixotropic parameters $a$ and $b$ to explore scenarios that involve changes in both the characteristic time $\lambda_{0}$ and the structure destruction parameter $\beta$. The density of both the fluid and the droplet is $\rho=1$. The maximum velocity reached by the fluid in the relaxed zone is $U=0.017$ in system units, and the associated sound velocity is $c=20>>U$. We run the simulation long enough to allow the droplet to cycle through the periodic channel several times and stabilize the effective viscosity. With the values described above, we obtain characteristic times of $\tau_{s}=0.8$, $\tau_{visc}=10$, and $T=200$, which comply with the relationship described in Equation 19. We first study the effect of varying the thixotropic time scale by covering four characteristic values $\lambda_{0}=[100,200,400]$, which are on the same time scale as the characteristic time of the fluid $T$. We analyze the viscosity evolution starting with both unformed ($f_{0}=0$) and formed ($f_{0}=1$) microstructure configurations. These results appear in Figure 9(a) and 9(b), respectively. In all the graphs, we normalize the viscosity using the limiting viscosity $\eta_{\infty}$ of the droplet, and we normalize the time axis using the expression $t^{\prime}=t\ \eta_{\infty}/(\rho L^{2})$. Figure 9(a) also shows a detail of the droplet when it is in the constricted zone and its respective relaxation in the channel zone with the largest area.
@@
Figure 9: Transient evolution of the viscosity of a thixotropic droplet passing through a chamber with diameter change. Comparison at different values of $\lambda_{0}$ at $\beta=1$ for cases starting with a) initially unformed microstructure $f_{0}=0$ and b) initially fully formed microstructure $f_{0}=1$.
For the case of initially unformed microstructure ($f_{0}=0$), it can be seen how the viscosity starts at its limiting value $\eta_{\infty}=10$ and evolves toward its maximum value defined by the thixotropic parameter $\alpha$. This evolution occurs more rapidly when the time scale is smaller, requiring fewer droplet cycles within the channel to reach a periodic stabilization of the viscosity. As $\lambda_{0}$ increases, more cycles are required, and the droplet may even fail to stabilize within the simulated cycles for the highest value of 400. For the case with initially formed microstructure ($f_{0}=1$), it can be seen how the viscosity starts at its maximum peak and decreases until it reaches a periodic stabilization. In this case, all studied values of $\lambda_{0}$ successfully achieve stabilization within the simulated cycles. It is clear that the viscosity decreases in the constriction zone and increases in the relaxed zone due to the abrupt change in the shear rate between these regions. We also investigate the effect of microstructure creation and destruction by varying the parameter $\beta$. We explore four values of $\beta=[0.1,1,10,100]$ for both unformed and formed microstructure. For these simulations, we set the characteristic time scale to $\lambda_{0}=200$, as this value aligns with the characteristic time of the fluid $T$. The results are shown in Figure 10(a) and 10(b), respectively. It can be seen that the direct effect of $\beta$ is to amplify the variations in shear rate and, consequently, the variations in the effective viscosity of the droplet. As the value of $\beta$ decreases, the influence of the shear rate on viscosity diminishes, leading to negligible viscosity variations for $\beta=0.1$ and to maximum variations for $\beta=100$.
@@
Figure 10: Transient evolution of the viscosity of a thixotropic droplet passing through a chamber with diameter change. Comparison at different values of $\beta$ at $\lambda_{0}=200$ for cases with a) initially unformed microstructure $f_{0}=0$ and b) initially fully formed microstructure $f_{0}=1$.
@@ 4.4 Droplet merging using micro-devices
Within microfluidic analysis, the process of droplet merging is of great interest. Various systems have been developed to measure the conditions and parameters involved in this phenomenon. Basu [50] developed a strategy known as Droplet Morphometry and Velocimetry (DMV), a video processing software that enables real-time observation of droplet behavior on small scales, particularly during merging events. Niu et al. [49] extended this research by employing microfluidic devices. To promote merging, they used a system of passive structures (pillars) inside a microfluidic channel, separated by specific distances, which caused droplet deceleration and eventual contact-induced coalescence. Using the SDPD multiphase methodology developed in our study, we numerically simulate this type of complex droplet interaction. Figure 11 shows the initial configuration for the merging process. The physical boundaries (in gray) represent the channel walls and pillar structures. The continuous phase (in white) flows through the channel, while two droplets are initially located at the entrance and exit of the system. We impose periodic boundary conditions in the $x$-direction so that as the final droplet reaches the left boundary of the domain, it reenters from the right as a new incoming droplet. We generate flow within the system by applying a body force. We analyze three particular cases. In the first case, both the penetrating droplet (orange in Figure 11) and the recipient droplet (purple) are Newtonian fluids with equal viscosity $\eta=0.1$. In the second case, the penetrating droplet is thixotropic with properties $\eta_{\infty}=0.1$, $\alpha=9$, $\lambda_{0}=100$, and $\beta=1$, while the recipient droplet remains Newtonian. In the third case, we reverse this configuration, considering the penetrating droplet as Newtonian and the recipient droplet as thixotropic with the same thixotropic parameters. In all cases, we consider the continuous phase as a Newtonian fluid with viscosity $\eta=0.1$. Figure 11 illustrates the merging dynamics. The first droplet enters the chamber and experiences significant resistance due to its interaction with the pillar geometry. Meanwhile, the second droplet advances and begins to penetrate the first droplet, which is partially trapped within the pillar structure. During this interaction, the second droplet transfers part of its kinetic energy to the first, facilitating the joint passage of both droplets through the pillar region. Upon exiting the interaction zone, it can be seen that the smaller droplet is almost fully embedded within the larger one. This interaction results from surface tension, which promotes coalescence and stabilizes the new droplet structure formed after fusion. Figure 11 also shows the final merged droplet for each studied case. It can be seen that in Case 1 (Newtonian/Newtonian), the penetrating droplet is nearly fully incorporated into the recipient droplet, resulting in a poorly defined internal structure with dispersed particles. In Case 2 (thixotropic/Newtonian), we observe a similar pattern, though with slightly more organized particles. Case 3 (Newtonian/thixotropic) exhibits a notably different result, with the formation of a flat and well-defined interface between the two droplets. This distinct boundary reflects the strong resistance imposed by the thixotropic recipient phase, which significantly slows down the penetration of the Newtonian droplet. As a consequence, the internal particle arrangement is more structured, clearly indicating the influence of the rheological properties of the recipient droplet on the merging dynamics.
@@
Figure 11: Initial state and merging process using microfluidic device. For Case 1, both droplets were considered to be Newtonian. In Case 2 the penetrating droplet (orange) is thixotropic and the recipient droplet (purple) is Newtonian. In Case 3 the penetrating droplet is Newtonian and the recipient droplet is thixotropic.
@@ 5 Conclusions
The validated thixotropic multiphase SDPD model was applied to a range of complex flow scenarios, demonstrating its potential for investigating biologically and industrially relevant systems. In the case of liquid–liquid phase separation (LLPS), the model captured how parameters such as protein concentration, capillary number, and microstructural relaxation influence droplet morphology—revealing conditions that favor the formation of stable, well-defined aggregates. Simulations of suspensions and emulsions showed that the model accurately reproduces the rheological evolution of complex fluids under confinement and shear, including effective viscosity trends and interfacial dynamics in two-phase systems. In microfluidic geometries, such as periodically constricted channels and droplet merging devices, the model captured key behaviors like localized viscosity variation, flow stabilization, and coalescence resistance due to internal structure. These results highlight the model’s capacity to simulate time-dependent rheology and interface behavior, making it a promising tool for the design and analysis of systems in soft matter physics, biomedical applications, and microfluidic device engineering.
@@ Appendix A Simple Newtonian Flows
We carried out the dynamic validation of the model considering several scenarios. The first part of the validation involves the development of a Poiseuille flow for a single-phase flow ($\alpha$) in three variations: simple profile, reverse Poiseuille and flow around a cylinder. For particle methods, as indicated by Backer et al. [61], the average velocity and shear stress in a Poiseuille flow are given by
[…]
@@
where $\rho$ is the density, $L$ the length between the parallel walls and $g_{x}$ the body force. For the simple and reverse case, we employ a two-dimensional square channel with walls separated by $L=10$, $\rho=1$ and $\eta=5$ in the units system. We select the equilibrium particle density $n_{eq}=25$ and we do not taken into account the effects of thermal fluctuations. We compare three velocities under laminar flow regime determine by the Reynolds number $Re=\rho v_{x}D/\eta$. We configure the body force with the required values to obtain the respective predefined velocities. In the case of reverse Poiseuille flow, the body force is positive in one half of the channel and negative in the opposite half, hence resulting in the associated representative velocity profile. We choose the speed of sound to be at least 20 times greater than the maximum velocity reached by the flow. We established a periodicity condition along the x-direction and we define a no-slip condition for walls by setting the related particles as totally stationary and without any contribution to the fluid. Figures 12(a), 12(b), 12(d) and 12(e) shows the velocity and shear profiles calculate by our model in comparison to the theoretical values obtain by applying Equation 20 for $Re=[0.1,0.5,1]$. Regarding the case of flow around a cylinder, we perform a validation based on the numerical results obtained by Ellero and Adams [40] using both SPH and immersed boundary method (IBM). We employ a channel characterized by a centrally located cylinder with radius $R_{c}=2$ and height $L=4R_{c}$ to study the flow characteristics. Considering the condition of periodicity on the x-direction, we set the width of the channel $L_{c}=6R_{c}$ such that the distance between repeated cylinders allows the fluid development. For this validation, the same physical parameters used in the first two cases are kept with the only exception of the viscosity whose new value is $\eta=1$ (in line with the approach of Ellero and Adams [40]). We configure the cylinder as well as the channel walls with a no-slip condition. Figures 12(c) and 12(f) shows the velocity profiles versus the normalized $x/R_{c}$ and $y/R_{c}$ axis for both for the results obtained with our approach and those getting by the IBM [40]. These curves are taken on a vertical plane at a distance of $x/R_{c}=3$ and a horizontal plane at a distance of $y/R_{c}=3.5$, respectively. As can be noted for the three analyses cases of single-phase flow, the comparative results between the theoretical values and the numerical approximations exhibit a close agreement showing the accuracy of the model in order to capture the characteristics of the flow under dynamic conditions.

### v2

@@ 4.3 Droplet dynamics in a periodically constricted channel
We now analyze the transition of a thixotropic droplet through a periodically constricted channel. Understanding the dynamics of thixotropic fluids in this setting is particularly relevant in systems where fluids experience recurrent deformation and relaxation, including blood flow in the microcirculation, polymeric suspensions in microfluidics, and drilling muds in porous media. The constricted geometry provides a controlled framework for examining the coupling between microstructural dynamics and macroscopic transport properties.
We consider a spherical droplet of size $R/\Delta x=10$ in a periodic channel of length $L_{cr}/\Delta x=160$, constituted by two sections of length $L/\Delta x=(L_{cr}/(\Delta x))/2=80$, with heights $H_{c}/\Delta x=11$ and $H_{r}/\Delta x=50$ for the constrained and relaxed sections, respectively (see Figure 9(a)). The maximum flow velocity in the channel is $U=0.017$. The relative lengths of the constriction and expansion determine whether the imposed shear aligns with the fluid’s thixotropic timescales of breakdown and recovery. Short constrictions may not fully degrade the structure, while longer ones drive the fluid toward its thinned state. Likewise, the expansion length controls how much recovery occurs before the next constriction. This coupling dictates local viscosity profiles and impacts transport efficiency and flow stability. We define the simulations parameters to ensure proper timescales hierarchy ${\tau_{\text{c}}}\ <<\ {\tau_{\text{vis}}}\ <<\tau_{\text{flow}},\lambda_{\text{thix}}$. Here, $\tau_{\text{c}}=0.8$ and the viscous relaxation time ${\tau_{\text{vis}}}=20$. The maximum velocity reached by the fluid in the relaxed zone leads to $\tau_{\text{flow}}=200$. To investigate the interplay between flowing and thixotropic time scales, we explore three specific conditions, $\lambda_{\text{thix}}=0.5\tau_{\text{flow}}$, $\lambda_{\text{thix}}=\tau_{\text{flow}}$, and $\lambda_{\text{thix}}=2.0\tau_{\text{flow}}$. This corresponds to a Reynolds number of $Re=0.17$ and a Thixotropic number $Th=0.005$. Additionally, we analyze the effects of microstructure destruction to formation rates by varying the parameter $\beta=\{0.1,1,10,100\}$. These experiments were conducted by by setting $\lambda_{\text{thix}}=\tau_{\text{flow}}$, resulting in $Re=0.17$ and $Th=\{0.0005,0.005,0.05,0.5\}$. For the different scenarios we consider droplets with initially unformed ($f_{0}=0$) and formed ($f_{0}=1$) microstructure. We consider the transporting fluid to be Newtonian with viscosity $\eta=1$, whereas the droplet has $\eta_{\infty}=10$ and $\alpha=4$ ($\eta_{\text{max}}=5\eta_{\infty}$). The density of both the fluid and the droplet is $\rho=1$.
@@
Figure 9: Transient evolution of the viscosity of a thixotropic droplet passing through a chamber with diameter change. Comparison at different values of $\lambda_{\text{thix}}$ at $\beta=1$ for cases starting with a) initially unformed microstructure $f_{0}=0$ and b) initially fully formed microstructure $f_{0}=1$.
In Figure 9, we provide the results for the droplet’s viscosity evolution for different thixotropic time scales, at a fixed destruction-to-formation rate $\beta=1$. For the case of initially unformed microstructure $f_{0}=0$ (see Figure 9(a)), it can be seen how the viscosity starts at its limiting value ($\eta/\eta_{\infty}=1$) and evolves toward its maximum. As expected, for shorter $\lambda_{\text{thix}}$, the stabilization of the viscosity towards $\eta_{\text{max}}$ occurs in a fewer constraint-to-relaxed cycles within the channel, whereas for the largest $\lambda_{\text{thix}}$, the full-microstructure formation is not even reached after seven cycles on the domain. In contrast, in droplets with initially formed microstructure $f_{0}=1$ (see Figure 9(b)), the viscosity profiles for the different $\lambda_{\text{thix}}$, readily converge within two cycles in the domain.
For lower value of $\beta=1$ explored, the destruction of the microstructure do not play a major role during the early stages of the droplet flow. However, as $\beta$ increases, it is expected that larger viscosity variation can emerge as the abrupt change in shear rate between regions occurs. In Figure 10, we compile the results of the viscosity variation for droplets with fixed thixotropic time scale to $\lambda_{\text{thix}}=\tau_{\text{flow}}$, for different values of $\beta$ and $f_{0}$. Consistently, we observe that the larger rates of microstructural destruction amplify the variations in shear rate, and consequently, the effective viscosity of the droplet.
In general, the observed interplay suggests that by tuning fluid properties together with channel dimensions, devices can be specifically designed to either minimize viscosity fluctuations for stable transport or exploit them for enhanced mixing, separation, or controlled delivery. For instance, lab-on-a-chip platforms can exploit viscosity variations to enhance mixing efficiency in low-Reynolds-number regimes, while separation devices can leverage controlled differences in viscosity to direct particles or cells along distinct trajectories. This ability to suppress or amplify viscosity fluctuations through the combined tuning of fluid properties and channel geometry underscores a powerful design principle for engineering application-specific fluid behaviors.
@@
Figure 10: Transient evolution of the viscosity of a thixotropic droplet passing through a chamber with diameter change. Comparison at different values of $\beta$ at $\lambda_{\text{thix}}=200$ for cases with a) initially unformed microstructure $f_{0}=0$ and b) initially fully formed microstructure $f_{0}=1$.
@@ 4.4 Droplet merging using micro-devices
Within microfluidic analysis, the conditions and parameters governing droplet merging of great interest niu2008pillar , basu2013droplet . Experimental techniques such as Droplet Morphometry and Velocimetry basu2013droplet (DMV) allow real-time observations of merging events on microscales resolutions. Moreover, microfluidic devices equipped with passive structures configurations niu2008pillar have been also successfully designed control contact-induced droplet coalescence. Here, we explore numerically such type of complex merging schemes when one of the fluids is thixotropic.
In Figure 11, we illustrate the merging domain configuration investigated. The physical boundaries (in gray) represent the channel walls and pillar structures. The continuous phase (in white) flows through the channel. We impose periodic boundary conditions in the flow direction, and apply a body force to induce the flow. We analyze three particular cases for which we set a sonic time scale $\tau_{\text{c}}=0.14$, interfacial tension time scale $\tau_{\sigma}=10$, viscous time scale $\tau_{\text{vis}}=7.84$, characteristic flow time scale $\tau_{\text{flow}}=490$ and thixotropic time $\lambda_{\text{thix}}=100$, following the hierarchy $\tau_{\text{c}}\ll\tau_{\text{vis}},\tau_{\sigma}\ll\tau_{\text{flow}},\lambda_{\text{thix}}$. In the first case, both the penetrating droplet (orange in Figure 11) and the recipient droplet (purple) are Newtonian fluids with equal viscosity $\eta=1$. In the second case, the penetrating droplet is thixotropic with properties $\eta_{\infty}=1$, $\alpha=9$ (this is $\eta_{\text{max}}=10\eta_{\infty}$), $\beta=1$, and $f_{0}=1$, while the recipient droplet remains Newtonian with $\eta=1$. The third case has the same configuration as the second case, with the only difference being that $\beta=500$. In all cases, we consider the continuous phase as a Newtonian fluid with viscosity $\eta=1$ and an input surface tension of $\sigma_{0}=2$. The height of the chamber in the central zone is equal to $H/\Delta x=220$ and the length $L/\Delta x=130$. The radius of the final droplet is calculated as $R/\Delta x=(\sqrt{N_{part}/d_{eq}})/\pi=20$. The maximum resultant velocity of the system is $U=0.09$ which results in a Capillary number of $Ca=0.045$, Reynolds number of $Re=2.4$ and a Thixotropic number of $Th=0.002$ and $Th=1$ for the two values of $\beta$.
In our simulation results, we observe that as the first droplet enters the chamber and experiences significant resistance due to its interaction with the pillar geometry. Meanwhile, the second droplet advances and begins to penetrate the first droplet, which is partially trapped within the pillar structure. During this interaction, the second droplet transfers part of its kinetic energy to the first, facilitating the joint passage of both droplets through the pillar region. Upon exiting the interaction zone, the smaller droplet is almost fully embedded within the larger one. This interaction results from surface tension, which promotes coalescence and stabilizes the new droplet structure formed after fusion.
Figure 11 compares the geometrical features of the final merged droplets for the three cases analyzed. In particular, we analyze the variations in the interfacial length $L_{\text{int}}/R$ between droplets ($R$ is the initial radius of the droplets) and the mean contact angle, given by $\theta_{\text{mean}}=(\theta_{\text{upper}}+\theta_{\text{lower}})/2$. For the Newtonian-Newtonian merging (case 1), the penetrating droplet retains a semicircular cap shape inside the recipient, with an interfacial length $L_{\text{int}}/R=4.7$, and mean angle $\theta_{\text{mean}}=31^{\circ}$. In the thixotropic-Newtonian case with $\text{Th}=0.002$ (case 2), the interfacial coverage is flatter, characterized by $L_{\text{int}}/R=4.3$ and $\theta_{\text{mean}}=44.1^{\circ}$. In the case 3, with $\text{Th}=1$, the interface adopts an intermediate configuration with $L_{\text{int}}/R=4.53$ and $\theta_{\text{mean}}=35.7^{\circ}$.
The differences between the three cases reflect how the hierarchy of time scales controls the dynamics of fusion. In all systems, $\tau_{\sigma}\ll\lambda_{\text{thix}}\ll\tau_{\text{flow}}$ holds, so that interface formation occurs much faster than microstructural reconstruction, which in turn occurs faster than global convective transport. In the case of low Th, the evolution of the microstructure is slow, retaining a higher viscosity during the merging and resulting in a flatter interface. Conversely, when Th increases, microstructural destruction occurs on a much smaller time scale, reducing the effective viscosity before contact and favoring deeper and more symmetrical penetration, similar to that observed in the Newtonian case. Taken together, these results show that the relationship between the scales $\lambda_{\text{thix}}$ and $\tau_{\text{flow}}$ determines the dominant regime of the process. When $\lambda_{\text{thix}}\ll\tau_{\text{flow}}$, the drop tends to behave like a more viscous fluid that is resistant to deformation, while, as $\lambda_{\text{thix}}$ increases or internal relaxation accelerates (higher Th), coalescence becomes faster and the interface adopts a greater curvature. These findings highlight two potential implications of thixotropic fluids in droplet-merging related process. First, the breadth and form of droplet fusion can be controlled by tuning factors the microstructural evolution of the fluid, offering an adjustable delay in the merging dynamics. Second, the difference between flat to circular penetration patterns illustrates the dynamic variation of the effective contact angle, which can be used in microfluidic applications that require controlled release, encapsulation, or mixing.
@@
Figure 11: Initial state and merging process using microfluidic device. For Case 1, both droplets were considered to be Newtonian. In Case 2 the recipient droplet (purple) is Newtonian and the penetrating droplet (orange) is thixotropic with $Th=0.002$. In Case 3 is the same that case 2 but $Th=1.0$.
@@ 5 Conclusions
The validated thixotropic multiphase SDPD model was applied to a range of complex flow scenarios, demonstrating its potential for investigating biologically and industrially relevant systems. In the case of liquid–liquid phase separation (LLPS), the model captured how parameters such as protein concentration, the capillary-driven coalescence, and microstructural relaxation influence droplet morphology—revealing conditions that favor the formation of stable, well-defined aggregates. Simulations of suspensions and emulsions showed that the model accurately reproduces the rheological evolution of complex fluids under confinement and shear, including effective viscosity trends and interfacial dynamics in two-phase systems.
In microfluidic geometries, such as periodically constricted channels and droplet merging devices, the model captured key behaviors like localized viscosity variation, flow stabilization, and coalescence resistance due to internal structure. Moreover, this results illustrate how tuning both the rheological parameters and the geometric design, it becomes possible to tailor viscosity profiles, shear stresses, and relaxation dynamics to specific functional needs. This coupling provides a framework for designing devices that exploit thixotropy, such as microfluidic systems for drug delivery, flow regulators in biomedical contexts, or extrusion processes in manufacturing.
Overall, our numerical results highlight the model’s capacity to simulate time-dependent rheology and interface behavior, making it a promising tool for the design and analysis of systems in soft matter physics, biomedical applications, and microfluidic device engineering.
Acknowledgments
The authors acknowledge funding by the Basque Government through the BERC 2022-2025 program and by the Ministry of Science, Innovation and Universities: BCAM Severo Ochoa accreditation CEX2021-001142-S / MICIN / AEI /10.13039/ 501100011033. The Spanish State Research Agency through the project PID2020-117080RB-C55 MICIU/AEI /10.13039 / 501100011033 funded by (AEI/FEDER, UE) with acronym COMPU - NANO - HYDRO.
CRediT authorship contribution statement
Andres Santiago Espinosa-Moreno: Investigation, Software, Validation, Writing – original draft, Writing – review & editing, Formal analysis. Nicolas Moreno: Conceptualization, Investigation, Supervision, Software, Methodology, Writing – original draft, Writing – review & editing, Formal analysis. Marco Ellero: Writing – review & editing, Supervision, Resources, Methodology, Funding acquisition, Formal analysis, Conceptualization.
@@ Appendix A Simple Newtonian Flows
We carried out the dynamic validation of the model considering several scenarios. The first part of the validation involves the development of a Poiseuille flow for a single-phase flow ($\alpha$) in three variations: simple profile, reverse Poiseuille and flow around a cylinder. For particle methods, as indicated by Backer et al. backer2005poiseuille , the average velocity and shear stress in a Poiseuille flow are given by
$v_{ave}$ $=<v_{x}>=\frac{1}{L}\int_{0}^{L}v_{x}(y)dy=\frac{\rho g_{x}L^{2}}{12\eta},$ (19)
@@
where $\rho$ is the density, $L$ the length between the parallel walls and $g_{x}$ the body force. For the simple and reverse case, we employ a two-dimensional square channel with walls separated by $L/\Delta x=50$, $\rho=1$ and $\eta=5$ in the units system. We select the equilibrium particle density $n_{eq}=25$ and we do not taken into account the effects of thermal fluctuations. We compare three velocities under laminar flow regime determine by the Reynolds number $Re=\rho v_{x}D/\eta$. We configure the body force with the required values to obtain the respective predefined velocities. In the case of reverse Poiseuille flow, the body force is positive in one half of the channel and negative in the opposite half, hence resulting in the associated representative velocity profile. We choose the speed of sound to be at least 20 times greater than the maximum velocity reached by the flow. We established a periodicity condition along the x-direction and we define a no-slip condition for walls by setting the related particles as totally stationary and without any contribution to the fluid. Figures 12(a), 12(b), 12(d) and 12(e) shows the velocity and shear profiles calculate by our model in comparison to the theoretical values obtain by applying Equation 19 for $Re=[0.1,0.5,1]$. Regarding the case of flow around a cylinder, we perform a validation based on the numerical results obtained by Ellero and Adams ellero2011sph using both SPH and immersed boundary method (IBM). We employ a channel characterized by a centrally located cylinder with radius $R_{c}/\Delta x=10$ and height $L=4R_{c}$ to study the flow characteristics. Considering the condition of periodicity on the x-direction, we set the width of the channel $L_{c}=6R_{c}$ such that the distance between repeated cylinders allows the fluid development. For this validation, the same physical parameters used in the first two cases are kept with the only exception of the viscosity whose new value is $\eta=1$ (in line with the approach of Ellero and Adams ellero2011sph ). We configure the cylinder as well as the channel walls with a no-slip condition. Figures 12(c) and 12(f) shows the velocity profiles versus the normalized $x/R_{c}$ and $y/R_{c}$ axis for both for the results obtained with our approach and those getting by the IBM ellero2011sph . These curves are taken on a vertical plane at a distance of $x/R_{c}=3$ and a horizontal plane at a distance of $y/R_{c}=3.5$, respectively. As can be noted for the three analyses cases of single-phase flow, the comparative results between the theoretical values and the numerical approximations exhibit a close agreement showing the accuracy of the model in order to capture the characteristics of the flow under dynamic conditions.